The area bounded by the curves y = x2 + 2 and y = 3x + 2 from intersection point to intersection point is rotated around x = 3.

Determine the formula for the the solid of revolution which is obtained from this rotation.


1
Expert's answer
2021-07-31T13:48:30-0400

The volume of the solid bounded by "y = x^2 + 2 \space and \space y = 3x + 2" about x-axis

"x^2 + 2 = 3x + 2 \\
x^2-3x=0\\
x=0,3\\
Volume = \int_0^3 \pi [(3x+2)^2-(x^2+2)^2]dx\\"


The volume of the solid bounded by "y = x^2 + 2 \space and \space y = 3x + 2" about x=3

"x^2 + 2 = 3x + 2 \\
x^2-3x=0\\
x=0,3\\
x= 0\implies y= 2\\
x= 3 \implies y= 11\\
Volume = \pi \int_2^{11} \pi [(R(x))^2-(r(x))^2]dx\\
y= 3x+2 \implies x= \frac{y-2}{3}\\
y=x^2+2 \implies x= \sqrt{y-2}\\
R(x) = 3- (\frac{y-2}{3})\\
r(x)= 3-\sqrt{y-2}\\
Volume = \pi \int_2^{11} \pi [(3- (\frac{y-2}{3}))^2-(3-\sqrt{y-2})^2]dx\\"